jake can complete a job in 5 hours while jerry can complete the same job in 3 hours. how long would it take to complete the job if jake starts the job two hours before jerry joins him
Answer by stanbon(75887)

jake can complete a job in 5 hours while jerry can complete the same job in 3 hours. how long would it take to complete the job if jake starts the job two hours before jerry joins him
----
Jake's rate: 1/5 job/hr
Jake's work time: x hrs
----------------------------
Jerry's rate: 1/3 job/hr
Jerry's work time: x-2 hrs
--------
Equation:
work + work = 1 job
----
x(1/5) + (x-2)(1/3) = 1
Multiply thru by 15 to get
3x + 5(x-2) = 15
8x - 10 = 15
8x = 25
x = 3 1/25 hr = 3 hr 2.4 minutes
===================================
Cheers,
Stan H.